OpenStudy (anonymous):

the school choir has 84 members. the ratio of girls to boys in the chor is 3.4 how many members are girls?

OpenStudy (anonymous):

3+4 = 7 84/7 = 12 Girls:Boys 3:4 Therefore, 3*12=36 girls There are 36 girls in the choir.